Jessica drove west at an average speed of 41 mph. Paul left from the same point in the opposite direction with an average speed of 46 mph. after Jessica had driven for 5 hours they were 389 miles apart. what is the number of hours Paul drove?
Answer by mananth(16949)

Jessica drove west at an average speed of 41 mph. Paul left from the same point in the opposite direction with an average speed of 46 mph. after Jessica had driven for 5 hours they were 389 miles apart. what is the number of hours Paul drove?
Jessica speed = 41 mph
Jessica travels 41*5 = 205 miles in 5 hours.
That is the time Paul starts
They are 389 miles away after x hours of paul starting
But Jessica had already traveled 205 niles
difference = 389-205 = 184 miles . they drove together
combined speed = 41+46 = 87 mph
distance = 184 miles
so time taken to be that far away = 184/87
= 2.11 hours
Paul drives 2.11 hours